OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = x2 – 2x + 1, what is f(g(x)) ?

OpenStudy (anonymous):

\[f(g(x))=f(x^2-2x+1)=4(x^2-2x+1)+1\] you can clean it up with algebra at the end if you like

OpenStudy (anonymous):

get \(f(g(x))=4x^2-8x+5\)

OpenStudy (anonymous):

thankyouuuu !!!!!

OpenStudy (anonymous):

yw